OpenStudy (anonymous):

x=30+2(s) where x+s=180. What is the value of x and s? Explain

OpenStudy (anonymous):

hmm lets seeeeeee x= 32? so what does s equal? like 180 - 32= s

OpenStudy (anonymous):

get it?

OpenStudy (anonymous):

and it would be 148 so 32 + 148 = 180

OpenStudy (anonymous):

Can you explain further?

OpenStudy (anonymous):

ohs i didn't see you replieddd >O<

OpenStudy (anonymous):

like it says :o.. x= 30 +2 right sos... it'd be 32 and 32-180 is? /o.o\ 148 get it? Q.Q

OpenStudy (anonymous):

Thankyou so much!!!!!!!!!!